OpenStudy (anonymous):

hehe em i have seen how wolframalhpa solve this this integration, but is it possible to solve it if i choose not to expand all this equation out? ¶ (3x^2+ 7x^3)^3 dx

OpenStudy (anonymous):

No

OpenStudy (anonymous):

so even those it is (ax+bx^2)^99, we have to expand them all??

OpenStudy (anonymous):

You could use the chain rule here. The chain rule goes like this: Suppose you have a 2 functions f(x) and g(x), and f(x) and g(x) is differential at any point c (f'(c) and g'(c) exists) Then: \[(f o g)(c) = f \prime(g(c)) * g \prime(c)\] In your example: Let \[f(x) = x ^{99}\] \[g(x) = ax + b ^{2}\] by applying chain rule: \[f(g(x) = (ax + b x^{2})^{99} = (ax + bx ^{2})^{98} * (a + 2bx)\]

OpenStudy (anonymous):

Use of integration by parts

OpenStudy (anonymous):

You would anyway have to do same no of steps

OpenStudy (anonymous):

oh sorry it says integration: Yeah you could do IBP

OpenStudy (anonymous):

how to do 'integration by part' ? and wat is IBP?

OpenStudy (anonymous):

IBP is integration by parts.

OpenStudy (anonymous):

How do u separate this equation? (ax+bx^2)^99

OpenStudy (anonymous):

separate them into ax and bx^2 then use the chain rule?

OpenStudy (anonymous):

This is pretty difficult you know, but alas i found the answer. Integration by parts goes like this: \[\int\limits f(x) g'(x)\, dx = f(x) g(x) - \int\limits f'(x) g(x)\, dx\] or simply, \[ \int\limits u\, dv=uv-\int\limits v\, du\! \] Now to answer your question, lets extract the x from the "inner equation" \[(3x^{2} + 7x^{3})^{3} = x^{6}(3+7x)^{3}\] Therfore, \[\int\limits_{}^{}(3x^{2}+7x^{3})^{3}dx=\int\limits_{}^{}x^{6}(3+7x)^{3}dx\] Using IBP, Let \[dv = (3+7x)^{3}dx\] \[v = \frac{(3 + 7x)^{4}}{28}\] \[u = x^{6}\]\[du = 6x^{5}dx\] Solving the equation above: \[∫x^{6}(3+7x)^{3}dx = \frac{(3+7x)^{4}}{28}* x^{6} - \int\limits_{}^{}\frac{(3+7x)^{4}}{28}*6x^{5}dx\] Notice that we still have a integration to solve at the right side. I'll tackle that on next post.

OpenStudy (anonymous):

So far we have solved the u * v part of the IBP but we are far from over. We still have an integration at the left. Again, lets do IBP, Let: \[dv _{1} = (3 + 7x)^{4}dx\]\[v _{1} = \frac{(3 + 7x)^{5}}{35}\]\[u _{1} = \frac{6}{28}x^{5}\]\[du _{1} = \frac{30}{28}x^{4}dx\] Solving again from IBP: \[−∫\frac{(3+7x)^{4}}{28}∗6x^{5}dx = - (\frac{(3+7x)^{5}}{35}*\frac{6}{28}x^{5} - \int\limits_{}^{}\frac{(3+7x)^{5}}{35}*\frac{30}{28}x^{4}dx\] Attaching the earlier result: \[∫x^{6}(3+7x)^{3}dx=\frac{(3+7x)^{4}}{28}∗x^{6}−\frac{(3+7x)^{5}}{35}∗\frac{6}{28}x^{5}+∫\frac{(3+7x)^{5}}{35}∗\frac{30}{28}x^{4}dx\] Notice the pattern here? If we repeat the IBP until the power of x is 0, we would have terms with powers of (3 + 7x) and x raise to something. Also the denominator = 7 * 7 * 5 * 4 or \[7^{2} * \frac{5!}{3!}\] the sign also alternates. So if we continue the trend: \[∫x^{6}(3+7x)^{3}dx = \frac{3!6!(3+7x)^{4}}{7*4!*6!}∗x^{6}−\frac{3!6!(3+7x)^{5}}{7^{2}*5!*5!}∗x^{5}+\frac{3!6!(3+7x)^{6}}{7^{3}*6!*4!}∗x^{4} - \] \[\frac{3!6!(3+7x)^{7}}{7^{4}*7!*3!}∗x^{3}+\frac{3!6!(3+7x)^{8}}{7^{5}*8!*2!}∗x^{2}-\frac{3!6!(3+7x)^{9}}{7^{6}*9!*2!}∗x^{2} + \frac{3!6!(3+7x)^{10}}{7^{7}*10!*1!}∗x^{1} - \] \[\frac{3!6!(3+7x)^{11}}{7^{8}*11!*0!}\]

OpenStudy (anonymous):

wow thats a long post XD